1.Calculate the pH of an aqueous solution of a 0.039 M HCI.

2. Calculate the pOH of an aqueous solution of .0.073 M LiOH.

3. The pOH of a solution of NaOH is 11.30. What is the [H*] for this solution?

4. The pH of a 0.02 M solution of an unknown weak acid is 3.7. what is the pKa of this acid?

5. What is the concentration of ammonium chloride in a solution if its pH is 4.80?

6. The pH of a solution is 4.80. What is the concentration of hydroxide ions in this solution?


1
Expert's answer
2021-10-22T03:51:47-0400

1. For a 0.025 M HCl solution the concentration of the hydrogen ions, [H+] , is 0.025 moles/liter (i.e. 0.025 M). pH is defined as −log[H+] . Substituting 0.025 into the equation for pH we find that the 0.025 M HCl solution has a pH of -log(0.039) = 1.41

pH= 1.41


2. pH = -log 0.073= 1.14

pH= 14+pOH

pOH = 14-1.14= 12.86